OpenStudy (anonymous):

145 is 16% of what number, how do I even start to solve this?

OpenStudy (anonymous):

? anyone

OpenStudy (anonymous):

@ducckyluver143 can you help me

OpenStudy (anonymous):

x of 16=145 x *16/100=145 then simplify by cross multiplying 16x=14500 then find value of x by dividing both sides by 16 x= 906.25

OpenStudy (anonymous):

thanks could you go more in depth, @Mariam.A

zepdrix (zepdrix):

It helps a little bit if you know what the key words actually mean. `is` means `=` `of` means `multiply` `per cent` means `per 100` `what number` is our unknown that we'll call `x`. 145 `is` `16` `%` `of` `what number` `145` `=` `16` `/100` `*` `x`

zepdrix (zepdrix):

\[\Large\rm 145=\frac{16}{100}\cdot x\]

zepdrix (zepdrix):

And then it requires a few algebra steps to solve for `x` from there, as Mariam showed.

OpenStudy (anonymous):

what are those few steps

OpenStudy (anonymous):

i also could use help on this question I don't know what to do

OpenStudy (anonymous):

Soo will rent a car for the weekend. She can choose one of two plans. The first plan has an initial fee of$44and costs an additional$0.14per mile driven. The second plan has an initial fee of$51and costs an additional$0.12per mile driven.

OpenStudy (anonymous):

for what amount of driving would the 2 plans cost the same? how would i calculate that

OpenStudy (anonymous):

and what is the cost when the 2 plans cost the same

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Nnesha could you help when you get a second

OpenStudy (anonymous):

ill help you! and sorry i was taking my unit test for my online school ;(

OpenStudy (anonymous):

@ducckyluver143 thanks man i'm really lost on this but i moved it to http://openstudy.com/study#/updates/542725c7e4b05a1848de120a